# Spectral Super-Resolution Neural Operator with Atmospheric Radiative Transfer Prior

### v1

@@ IV-D Ablation Study
IV-E Continuous and Zero-shot Spectral Reconstruction
To further evaluate the generalization and flexibility of our method, we conduct experiments on two challenging scenarios: continuous spectral reconstruction and zero-shot spectral reconstruction.
In the continuous reconstruction setting, we train the model using only a small number of uniformly downsampled spectral bands and test it on the full spectral resolution. This simulates an interpolation task in the spectral domain, where the model needs to infer missing spectral information between known bands. In the zero-shot setting, we train the model using only the visible spectrum range but test the model on the complete spectral range from 400 nm to 2500nm, which evaluates the ability to reconstruct the entire hyperspectral signal, including the never-seen infrared region. This corresponds to a spectral extrapolation task, which is highly challenging due to the lack of direct observation in the infrared spectral range.
In our experiments, we apply $2\times$ downsampling for continuous spectral reconstruction and restrict the spectrum to wavelengths below 1000 nm for zero-shot reconstruction. We compare our method against two other arbitrary-scale methods: NeSR and NeSSR. Table V summarizes the quantitative results. Our method achieves best performance in both settings, indicating strong generalization to unseen or sparsely-sampled spectral distributions.
The good performance of SSRNO stems from two factors: the neural operator backbone ensuring resolution-invariant spectral interpolation, and the ART prior guiding plausible extrapolation from visible to infrared bands. These results highlight SSRNO’s robustness and generalization ability, showing that it can both interpolate missing spectra and extrapolate beyond observed bands, making it suitable for practical hyperspectral applications.
@@ V Conclusion
In this article, we propose a novel physics-informed approach for remote sensing image spectral reconstruction. By embedding physical priors into a data-driven framework through the GMP method, our method effectively balances interpretability and adaptability. Experiments across diverse geographical scenes indicate better reconstruction accuracy and robustness, especially in the challenging continuous and zero-shot spectral reconstruction scenarios. This work highlights the potential of integrating domain knowledge with deep learning to advance spectral super-resolution in real-world remote sensing applications.

### v2

@@ IV-E Ablation Study
In this section, we conduct ablation studies to further investigate the effectiveness of the proposed method. The key innovation of our method lies in the ACP, which integrates physical priors into a data-driven pipeline during the initial upsampling and enforces hard constraints during final refinement. To validate its effectiveness, we perform ablation studies on the upsampling stage and the refinement stage. For the upsampling stage, we use a zero matrix as the angular guidance $Z$ to obtain the upsampled HSI, rather than using the radiative prior, which is also equivalent to directly employing the Moore-Penrose inverse. To compare hard and soft constraints, we directly remove the final refinement stage and add an extra error term on the loss function as the regularization to enforce low reconstruction error, which is
$\mathcal{L}_{r}(\hat{X},X)=\mathcal{L}(\hat{X},X)+\alpha L_{1}(S_{\boldsymbol{w}}\hat{X},S_{\boldsymbol{w}}X),$ (29)
where $\alpha$ is the trade-off parameter, and we set $\alpha=0.5$ in the ablation experiment. The results are listed in Table V.
The results indicate that both the radiative prior and the final refinement play critical roles in the overall performance. Removing the radiative prior in the upsampling stage reduces PSNR and increases SAM, showing that physically informed guidance is crucial for producing an initial estimate close to the target. However, eliminating the refinement stage and replacing it with soft regularization degrades both PSNR and SAM more significantly, suggesting that hard constraints in the output space are more effective than radiative prior in preserving spectral fidelity. Overall, these findings highlight that the ACP method plays a crucial role in ensuring accurate and physically consistent spectral reconstruction.
@@ V Conclusion
In this article, we propose a novel physics-informed approach for continuous and extrapolative spectral reconstruction. By continuously reformulating the problem and embedding physical priors within a data-driven framework, our method yields more realistic reconstructions and effectively balances interpretability and adaptability. Experiments across diverse real-world scenes and problem settings indicate better reconstruction accuracy and robustness, especially in the continuous setting. This work highlights the potential of continuous modeling of spectra and integrating domain knowledge with deep learning to advance spectral super-resolution in practical applications.